# MatAnyone: Stable Video Matting with Consistent Memory Propagation

### v1

@@ 2 Related Work
Video Matting. Due to the intrinsic ambiguity in the auxiliary-free setting [57, 39, 33, 24, 27, 61], such tasks generally are object-specific. Among them, human video matting [43, 61, 24, 27] without auxiliary inputs is popular due to its wide applications. Challenging as the auxiliary-free setting, being in the video domain brings in additional difficulties in temporal coherency. MODNet [24] extends its portrait matting setting to video domain with a flickering reduction trick (non-learning) within a local sequence. RVM [33] steps further to design for videos specifically with ConvGRU [1] as its recurrent architecture. Robust as RVM, it is still easy to be confused by humans in the background. With the success of promptable segmentation [25, 40, 62, 58], obtaining segmentation mask for a target human object only requires minimal human efforts. Recent mask-guided image [56, 55, 3, 29] and video matting [21, 31, 22, 28] thus leverages this convenience for a more robust performance. Adam [31] propagates the first-frame segmentation mask across all frames while FTP-VM [21] propagates the first-frame trimap. Taking the propagated mask as a rough result, their decoder serves for matting details refinement. MaGGIe [22] enjoys a stronger prior by taking the segmentation mask across all frames instead of the first one. Taking all the segmentation masks at a time, the network is able to perform bidirectional temporal fusion for coherency. To mitigate the poor generalizability of synthetic video matting data, a common practice is to simultaneously train with real segmentation data for semantics supervision [33, 31, 21].
@@
Video Consistency in Low-level Vision. To enhance temporal consistency across adjacent frames, the recurrent frame fusion [59, 47] and optical flow-guided propagation [4, 5, 6] are commonly utilized in the video restoration networks. Recent methods also employ temporal layers such as 3D convolution [2, 48] and temporal attention [2, 7, 60, 49] during training, while other training-free methods resorts to cross-frame attention [50, 53] and flow-guided attention [15, 13] in the pretrained models. In this work, we find that the memory-based paradigm is effective enough to maintain video consistency for video matting.
@@ 3 Methodology
We also propose two practical inference strategies that allow for flexible application, 1) a recurrent refinement approach based on the memory-driven paradigm, and 2) an auxiliary-free variant that eliminates the need for a target segmentation mask in the first frame (Sec. 3.3).
@@ 3.2 Core-area Supervision via Segmentation
MODNet [24] RVM [33] RVM-Large [33] Ours-AF AdaM [31] FTP-VM [20] MaGGIe [22] Ours
@@
3.3 Practical Inference Strategies
Recurrent Refinement. The first-frame matte is predicted from the given first-frame segmentation mask, and its quality will affect the matte prediction for the subsequent frames. The sequential prediction in the memory-based paradigm enables recurrent refinement during inference. Leveraging this mechanism, we introduce an optional first-frame warm-up module for inference. Specifically, we repeat the first frame $n$ times, treating each repetition as the initial frame, and use only the $n^{th}$ alpha output as the first frame to initialize the alpha memory bank. This (1) enhances robustness against the given segmentation mask and (2) refines matting details in the first frame to achieve image-matting quality (see Fig. 6 and Fig. 13).
Auxiliary-free Variant. To enable comparison with an arbitrary auxiliary-free video matting approach, we design an auxiliary-free version by removing the segmentation prior from the initial frame. Instead, we use the first-frame alpha matte generated by an auxiliary-free method of interest, such as RVM, and binarize it as the given mask for our setting. Table 1 presents a comparison between RVM and our auxiliary-free variant (Ours-AF) on synthetic benchmarks.
@@ 5.1.1 Quantitative Evaluations
Synthetic Benchmarks. For a comprehensive evaluation on synthetic benchmarks, we employ MAD (mean absolute difference) and MSE (mean squared error) for semantic accuracy, Grad (spatial gradient) [41] for detail extraction, Conn (connectivity) [41] for perceptual quality, and dtSSD [14] for temporal coherence. In Table 1, our method achieves the best MAD and dtSSD across all datasets at both high and low resolutions, demonstrating exceptional spatial accuracy for alpha mattes and remarkable temporal stability. Our auxiliary version, which shares the same first-frame prediction as RVM, outperforms RVM in both dtSSD and Conn metrics across all datasets, highlighting the advantages of our design in stability and visual quality.

### v2

@@ 2 Related Work
Video Matting. Due to the intrinsic ambiguity in the auxiliary-free setting [57, 39, 33, 24, 27, 62], such tasks generally are object-specific. Among them, human video matting [43, 62, 24, 27] without auxiliary inputs is popular due to its wide applications. Challenging as the auxiliary-free setting, being in the video domain brings in additional difficulties in temporal coherency. MODNet [24] extends its portrait matting setting to video domain with a flickering reduction trick (non-learning) within a local sequence. RVM [33] steps further to design for videos specifically with ConvGRU [1] as its recurrent architecture. Robust as RVM, it is still easy to be confused by humans in the background. With the success of promptable segmentation [25, 40, 63, 58], obtaining segmentation mask for a target human object only requires minimal human efforts. Recent mask-guided image [56, 55, 3, 29] and video matting [21, 31, 22, 28] thus leverage this convenience for a more robust performance. Adam [31] propagates the first-frame segmentation mask across all frames while FTP-VM [21] propagates the first-frame trimap. Taking the propagated mask as a rough result, their decoder serves for matting details refinement. MaGGIe [22] enjoys a stronger prior by taking the segmentation mask across all frames instead of the first one. Taking all the segmentation masks at a time, the network is able to perform bidirectional temporal fusion for coherency. To mitigate the poor generalizability of synthetic video matting data, a common practice is to simultaneously train with real segmentation data for semantic supervision [33, 31, 21].
@@
Video Consistency in Low-level Vision. To enhance temporal consistency across adjacent frames, the recurrent frame fusion [59, 47] and optical flow-guided propagation [4, 5, 6, 60] are commonly utilized in the video restoration networks. Recent methods also employ temporal layers such as 3D convolution [2, 48] and temporal attention [2, 7, 61, 49] during training, while other training-free methods resort to cross-frame attention [50, 53] and flow-guided attention [15, 13] in the pretrained models. In this work, we find that the memory-based paradigm is effective enough to maintain video consistency for video matting.
@@ 3 Methodology
We also propose a practical inference strategy that allow for flexible application: a recurrent refinement approach applied to the first frame, based on the memory-driven paradigm, enhancing robustness to the given mask and refining matting details (Sec. 3.3).
@@ 3.2 Core-area Supervision via Segmentation
MODNet [24] RVM [33] RVM-Large [33] AdaM [31] FTP-VM [20] MaGGIe [22] Ours
@@
3.3 Recurrent Refinement During Inference
The first-frame matte is predicted from the given first-frame segmentation mask, and its quality will affect the matte prediction for the subsequent frames. The sequential prediction in the memory-based paradigm enables recurrent refinement during inference. Leveraging this mechanism, we introduce an optional first-frame warm-up module for inference. Specifically, we repeat the first frame $n$ times, treating each repetition as the initial frame, and use only the $n^{th}$ alpha output as the first frame to initialize the alpha memory bank. This (1) enhances robustness against the given segmentation mask and (2) refines matting details in the first frame to achieve image-matting quality (see Fig. 6 and Fig. 13 in the appendix).
@@ 5.1.1 Quantitative Evaluations
Synthetic Benchmarks. For a comprehensive evaluation on synthetic benchmarks, we employ MAD (mean absolute difference) and MSE (mean squared error) for semantic accuracy, Grad (spatial gradient) [41] for detail extraction, Conn (connectivity) [41] for perceptual quality, and dtSSD [14] for temporal coherence. In Table 1, our method achieves the best MAD and dtSSD across all datasets at both high and low resolutions, demonstrating exceptional spatial accuracy for alpha mattes and remarkable temporal stability. Apart from accuracy and stability, our method achieves the best Conn on both benchmarks, indicating its superior visual quality (Fig. 4 and Sec. J.5 in the appendix).
@@ 6 Conclusion
Acknowledgement. This study is supported under the RIE2020 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s).